Ticket Number: HEA-NOC/20041017-3 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: Technico
Ticket Opened: 20041017 10:30 UTC+1 Problem Start: 20041017 09:05 UTC+1
Ticket Update: 20041018 16:13 UTC+1 Problem End: 20041018 12:00 UTC+1
Auto Close: 1 day
The Hermes access router is not responding to probes.
All traffic to/from the Hermes schools in Portmarnock
20041017 10:28 BB Investigating
20041017 10:52 BB The radio link to Hermes appears normal. Possibly the
power to the client access router has failed.
20041018 09:33 SY Contacted Hermes technical people, left a message for
Noel Connelly awaiting his reply.
20041018 16:10 BB The power supply for Technico's radio equipment based in
the DCU POP had failed. Technico replaced the power supply, and traffic is
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20041017-3